# Periodic phase slips and frequency comb generation at tunable microwave frequencies in superconducting diabolo structures

### v1

@@ Abstract
[…]
@@ 1 Introduction
However, while phase slip phenomena in one dimension are relatively well understood16, analytical descriptions in two and three dimensions remain challenging due to the complexity of the governing equations. In particular, the interplay between topological defects, such as vortices and phase slips, in 3D nanostructures is not fully understood.
In this work, we conduct full 3D time-dependent Ginzburg-Landau simulations to investigate phase slip phenomena in Nb diabolo structures — hollow tubes featuring a central constriction (Figure 1a) — under both DC and combined DC+AC current drives. We demonstrate that the intrinsic frequency of phase slip oscillations can be harnessed to generate microwave frequency combs, potentially advancing high-frequency quantum applications. Furthermore, we demonstrate the coexistence of phase slips with vortices and show that at sufficiently high currents, vortices are absorbed by the phase slip region, offering insights into controlling topological defects in superconducting systems.
@@ 2 Results and discussion
We have numerically simulated several tubular structures of different geometrical parameters and report here on a superconducting Nb diabolo structure (Figure 1a). It consists of an Nb shell with a thickness of $d=27$ nm and length $l=1.64~{}\mu$m. The radius varies from $r_{\rm o}=890$ nm at the ends to $r_{\rm i}=57$ nm at the central constriction. The penetration depth and coherence length are taken as $\lambda=273$ nm and $\xi=58$ nm, respectively. Since $d\ll\lambda$, self-field effects are neglected. To study the dynamic response, we numerically solve the time-dependent Ginzburg–Landau (TDGL) equations using a 3D finite-element formulation17 implemented in COMSOL Multiphysics18. We adopt the Coulomb gauge19, imposing $\nabla\cdot\mathbf{A}=0$ on the vector potential $\mathbf{A}$.
@@
Here, lengths are normalized to $\lambda$, and time to $\xi^{2}/D$, with $D$ being the diffusion coefficient. The last term in Eq. 1 is negligible in the limit of small self-fields. The evolution of $\psi$ is described by the TDGL equation:
@@
Additionally, to ensure uniqueness, we impose the condition:
@@
Figure 2b presents the instantaneous voltage $U$ as a function of time at an applied current of $j_{\rm DC}=0.36~{}\rm{MA~{}cm^{-2}}$. To understand this regime, we analyze the spatial distribution of the order parameter’s phase and magnitude (Figure 2c). At the voltage peak (indicated by a square symbol), the order parameter is fully suppressed at the constriction. The phase profile is rotationally symmetric and varies along the arc length of the diabolo. Figure 2d illustrates the phase variation (left) and order parameter (right) along the arc length for three different points in time. The phase undergoes a $2\pi$ jump at the constriction during each phase slip. As described by Michotte et al.20, the periodic phase slip state is stable when the relaxation time of the order parameter magnitude is shorter than that of the phase, i.e., $\tau_{|\psi|}<\tau_{\phi}$.
The phase slip regime persists in the presence of vortices. Figure 3 presents a DC current sweep performed at an applied field of 12.5 mT, where the diabolo hosts two vortex rows (see Figure 1c). At low DC currents, vortices and anti-vortices begin to move under the influence of the Lorentz force. Due to the opposite circulation of their screening currents, their motion under the transport current occurs in opposite directions—either clockwise or counter-clockwise. This motion results in a finite, time-constant voltage $U$. As the DC current increases, the diabolo transitions into the phase slip regime. Notably, the frequency of phase slips scales with the DC current (Figure 3b). Additionally, higher-order harmonics appear in the voltage spectrum, which are attributed to the system’s intrinsic non-linearity. The time-averaged voltage $\langle U\rangle$ and the corresponding differential voltage $d\langle U\rangle/dj$ are shown in Figure 3c and 3d, respectively. Both reveal distinct jumps at $j_{\rm DC}=0.6~{}\mathrm{MA~{}cm^{-2}}$ and $j_{\rm DC}=0.66~{}\mathrm{MA~{}cm^{-2}}$. During these transitions, the vortex and anti-vortex rows are absorbed into the constriction by the phase slip. Depending on $j_{\rm DC}$, this process leads to periodic vortex–anti-vortex pair nucleation and annihilation within the constriction (Figure 3e).
@@
The phase slips and their DC-current-controlled GHz oscillations open up an intriguing functionality. This is demonstrated by introducing an additional AC current modulation $j_{\mathrm{AC}}=j_{0}\sin(2\pi f_{\mathrm{AC}}t)$ to the DC current $j_{\mathrm{DC}}$ when the sample is in the phase slip regime (Figure 4a). Here, $j_{0}$ is the AC amplitude, which is fixed at $j_{0}=0.036~{}\mathrm{MA~{}cm^{-2}}$, and $f_{\mathrm{AC}}=15$ GHz is the modulation frequency. We start from an initial state with $B_{\mathrm{z}}=12.5$ mT and $j_{\mathrm{DC}}=0.36~{}\mathrm{MA~{}cm^{-2}}$, ensuring a stable phase slip regime (Figure 3). The AC amplitude $j_{0}$ is gradually ramped from 0 to $0.036~{}\mathrm{MA~{}cm^{-2}}$. Subsequently, the AC amplitude and frequency are held constant while sweeping over the DC amplitude. The resulting $j_{\rm DC}$-dependent spectra (Figure 4b) possess several noteworthy features. Instead of a single fundamental peak and its harmonics, the introduction of the AC signal leads to a range of harmonic peaks. This frequency mixing arises from the nonlinearity of the superconducting condensate, resulting in frequencies of the form $f=nf_{\mathrm{ps}}+mf_{\mathrm{AC}}$, where $n$ and $m$ are integers, and $f_{\mathrm{ps}}$ is the fundamental phase slip frequency driven by the DC current. This mechanism realizes a frequency comb with $n=1$ consisting of hierarchies of equidistant frequency peaks for a fixed $j_{\rm DC}$. Additionally, a second set of branches with a steeper slope is observed for small $j_{\rm DC}$ which corresponds to the $n=2$ harmonics. Further simulations indicate that the width of the frequency comb scales with the AC amplitude, while the comb spacing is determined by the modulation frequency $f_{\rm AC}$ (see Supplementary Information). A distinct jump in the spectrum is observed, which is reflected in both the time-averaged voltage and the differential voltage (Figure 4c). This discontinuity corresponds to the entrance of a single vortex and anti-vortex pair into the constriction.
@@
Prior to the jump, the phase evolution resembles that shown in Figure 2, where the phase distribution within the constriction exhibits rotational symmetry. At the onset of the jump, an isolated vortex first enters the constriction and moves at high velocities. Subsequently, at $t\approx 420$ ps, an anti-vortex enters the constriction. This leads to the disappearance of the higher-order frequency branches in the spectrum. At higher DC currents shown in Figure 4 b, no additional vortices are observed to enter the constriction. Moreover, upon subsequently reducing the DC current, the system does not exhibit a reappearance of the jump, featuring the presence of hysteresis in the current-voltage characteristic.
@@
Figure 4: DC current sweep of the diabolo under constant applied field and AC current. (a) schematic of the simulation, highlighting that the injected DC transport current is now modulated with an AC current at constant amplitude and frequency. (b) Resulting Fourier transformed voltage as a function of DC current amplitude. The spectrum shows non-linear frequency mixing of the phase slip frequency $f_{\rm ps}$ with the ac current frequency $f_{AC}$. The right side shows a linecut of the spectrum taken at $J_{DC}=0.39~{}\mathrm{MA~{}cm^{-2}}$ (dashed line). (c) Time averaged voltage (top) and differential time averaged voltage (bottom) as a function of DC current.
@@
We have investigated the phase slips in the tubular 3D nanostructures by employing the TDGL equations, which are valid for gapless superconductors near the critical temperature 21. Previous studies have shown that the onset current for phase slips depends on the relaxation times of both the order parameter’s magnitude and phase22, 20. The magnitude relaxation is influenced by the finite inelastic scattering time, which could be incorporated using the generalized TDGL formalism23. Although this refinement might quantitatively affect the results, the standard TDGL framework has been extensively applied to study phase slips and reliably captured the essential behavior qualitatively 24, 25.
@@ 2.1 Conclusion
Our findings open possibilities for developing 3D integrated, on-chip frequency comb generators by conformal coating of a pillar-like nanotemplate with a single superconducting shell. The diabolo structure as described here avoids an SIS junction and could also be realized by using a direct write method based on a focused ion beam 26, 27. Our design offers a promising advancement for quantum nanotechnologies and high-frequency applications.

### v2

@@ Abstract
Institute of Electrical and Micro Engineering, School of Engineering, École Polytechnique Fédérale de Lausanne (EPFL), Lausanne, 1015, Vaud, Switzerland
@@ 1 Introduction
However, while phase slip phenomena in one dimension are relatively well understood16, 17, analytical descriptions in two and three dimensions remain challenging due to the complexity of the governing equations. In particular, the interplay between topological defects, such as vortices and phase slips18, 13, in 3D nanostructures is not fully understood.
In this work, we conduct full 3D time-dependent Ginzburg-Landau simulations to investigate phase slip phenomena in Nb diabolo structures — hollow tubes featuring a central constriction (Figure 1a) — under both DC and combined DC+AC current drives. We demonstrate that the intrinsic frequency of phase slip oscillations, given by the Josephson relation 19, can be harnessed to generate microwave frequency combs, potentially advancing high-frequency quantum applications. Furthermore, we demonstrate the coexistence of phase slips with vortices and show that at sufficiently high currents, vortices are absorbed by the phase slip region, offering insights into controlling topological defects in superconducting systems.
@@ 2 Results and discussion
[…]
@@
Here, lengths are normalized to $\lambda$, and time to $\xi^{2}/D$, with $D$ being the diffusion coefficient. The last term in Eq. 1 is negligible in the limit of small self-fields. The evolution of $\psi$ is described by the TDGL equation25, 21:
@@
With $j_{\rm tr}$ the applied transport current. Additionally, to ensure uniqueness, we impose the condition:
@@
Figure 2b presents the instantaneous voltage $U$ determined as the difference in scalar potential between the top and bottom leads as a function of time at an applied current of $j_{\rm DC}=0.36~{}\rm{MA~{}cm^{-2}}$. To understand this regime, we analyze the spatial distribution of the order parameter’s phase and magnitude (Figure 2c). At the voltage peak (indicated by a square symbol), the order parameter is fully suppressed at the constriction. The phase profile is rotationally symmetric and varies along the arc length of the diabolo. Figure 2d illustrates the phase variation (left) and order parameter (right) along the arc length for three different points in time. The phase undergoes a $2\pi$ jump at the constriction during each phase slip. As described by Michotte et al.26, the periodic phase slip state is stable when the relaxation time of the order parameter phase is shorter than that of the magnitude, i.e., $\tau_{\phi}<\tau_{|\psi|}$.
The phase slip regime persists in the presence of vortices. Figure 3 presents a DC current sweep performed at an applied field of 12.5 mT, where the diabolo hosts two vortex rows (see Figure 1c). At low DC currents, vortices and anti-vortices begin to move under the influence of the Lorentz force. Due to the opposite circulation of their screening currents, their motion under the transport current occurs in opposite directions—either clockwise or counter-clockwise. This motion results in a finite, time-constant voltage $U$. As the DC current increases, the diabolo transitions into the phase slip regime. Notably, the frequency of phase slips scales with the DC current (Figure 3b). This frequency is determined by the Josephson relation, which relates the phase slip frequency $f_{\rm ps}$ to the average voltage drop over the phase slip $\bar{U}=\frac{h}{2e}f_{\rm ps}$19. Additionally, higher-order harmonics appear in the voltage spectrum, which are attributed to the system’s intrinsic non-linearity. The time-averaged voltage $\langle U\rangle$ and the corresponding differential voltage $d\langle U\rangle/dj$ are shown in Figure 3c and 3d, respectively. Both reveal distinct jumps at $j_{\rm DC}=0.6~{}\mathrm{MA~{}cm^{-2}}$ and $j_{\rm DC}=0.66~{}\mathrm{MA~{}cm^{-2}}$. During these transitions, the vortex and anti-vortex rows are absorbed into the constriction by the phase slip. Depending on $j_{\rm DC}$, this process leads to periodic vortex–anti-vortex pair nucleation and annihilation within the constriction (Figure 3e).
@@
The phase slips and their DC-current-controlled GHz oscillations open up an intriguing functionality. This is demonstrated by introducing an additional AC current modulation $j_{\mathrm{AC}}=j_{0}\sin(2\pi f_{\mathrm{AC}}t)$ to the DC current $j_{\mathrm{DC}}$ when the sample is in the phase slip regime (Figure 4a). Here, $j_{0}$ is the AC amplitude, which is fixed at $j_{0}=0.036~{}\mathrm{MA~{}cm^{-2}}$, and $f_{\mathrm{AC}}=15$ GHz is the modulation frequency. We start from an initial state with $B_{\mathrm{z}}=12.5$ mT and $j_{\mathrm{DC}}=0.36~{}\mathrm{MA~{}cm^{-2}}$, ensuring a stable phase slip regime (Figure 3). The AC amplitude $j_{0}$ is gradually ramped from 0 to $0.036~{}\mathrm{MA~{}cm^{-2}}$. Subsequently, the AC amplitude and frequency are held constant while sweeping over the DC amplitude. The resulting $j_{\rm DC}$-dependent spectra (Figure 4b) possess several noteworthy features. Instead of a single fundamental peak and its harmonics, the introduction of the AC signal leads to a range of harmonic peaks. This frequency mixing arises from the nonlinearity of the superconducting condensate, resulting in frequencies of the form $f=nf_{\mathrm{ps}}+mf_{\mathrm{AC}}$, where $n$ and $m$ are integers, and $f_{\mathrm{ps}}$ is the fundamental phase slip frequency driven by the DC current. This mechanism realizes a frequency comb with $n=1$ consisting of hierarchies of equidistant frequency peaks for a fixed $j_{\rm DC}$. Additionally, a second set of branches with a steeper slope is observed for small $j_{\rm DC}$ which corresponds to the $n=2$ harmonics. Further simulations indicate that the width of the frequency comb scales with the AC amplitude, while the comb spacing is determined by the modulation frequency $f_{\rm AC}$ (Supporting Information). A distinct jump in the spectrum is observed, which is reflected in both the time-averaged voltage and the differential voltage (Figure 4c). This discontinuity corresponds to the entrance of a single vortex and anti-vortex pair into the constriction.
@@
Prior to the jump, the phase evolution resembles that shown in Figure 2, where the phase distribution within the constriction exhibits rotational symmetry. At the onset of the jump, an isolated vortex first enters the constriction and moves at high velocities. Subsequently, at $t\approx 420$ ps, an anti-vortex enters the constriction. This leads to the disappearance of the higher-order frequency branches in the spectrum. At the higher DC currents shown in Figure 4 b, no additional vortices are observed to enter the constriction. Moreover, upon subsequently reducing the DC current, the system does not exhibit a reappearance of the jump, featuring the presence of hysteresis in the current-voltage characteristic.
@@
Figure 4: DC current sweep of the diabolo under constant applied field and AC current. (a) schematic of the simulation, highlighting that the injected DC transport current is now modulated with an AC current at constant amplitude and frequency. (b) Resulting Fourier transformed voltage $\tilde{U}$ as a function of DC current amplitude. The spectrum shows non-linear frequency mixing of the phase slip frequency $f_{\rm ps}$ with the ac current frequency $f_{AC}$. The right side shows a linecut of the spectrum taken at $j_{DC}=0.39~{}\mathrm{MA~{}cm^{-2}}$ (dashed line). (c) Time averaged voltage (top) and differential time averaged voltage (bottom) as a function of DC current.
@@
We have investigated the phase slips in the tubular 3D nanostructures by employing the TDGL equations, which are valid for gapless superconductors near the critical temperature 25. Previous studies have shown that the onset current for phase slips depends on the relaxation times of both the order parameter’s magnitude and phase27, 26. The magnitude relaxation is influenced by the finite inelastic scattering time, which could be incorporated using the generalized TDGL formalism28. Although this refinement might quantitatively affect the results, the standard TDGL framework has been extensively applied to study phase slips in one- and two-dimensional systems and reliably captured the essential behavior qualitatively 29, 8, 30.
@@ 2.1 Conclusion
Our findings open possibilities for developing 3D integrated, on-chip frequency comb generators by conformal coating of a pillar-like nanotemplate with a single superconducting shell. The diabolo structure as described here avoids an SIS junction and could also be realized by using a direct write method based on a focused ion beam 31, 32. Our design offers a promising advancement for quantum nanotechnologies and high-frequency applications.
@@
{suppinfo}
Additional details on numerical procedure, simulations on wider constrictions, frequency comb dependence on AC and DC current strength.